OpenStudy (anonymous):

Math help Write 3^x =12 in logarithmic form Write in 7=x in exponential form

OpenStudy (anonymous):

Logs simply answer the question "what power must 3 be raised to to get 12?" log 3^x =12

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

OK if y= b^x then log to the base b of y = x

OpenStudy (anonymous):

Example: if 100= 10^2 then log to trhe base 10 of 100 = 2

OpenStudy (anonymous):

generally what happens is you take the log of both sides: log (3^x) = log(12) then due to another rule: x. log3 = log12 and then x= log 12/log3 x= 2.26

OpenStudy (anonymous):

I am sure the real mathematicians in here (of which there are many) would explain it far better than me.

OpenStudy (anonymous):

However for the moment, I calculate 3^ 2.26=11.97551049 which is approximately 12

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You are welcome @wagnermar, actually its lovely to get revision of logs!

OpenStudy (anonymous):

P.S. if you see an equation of the form 3^x = 12 (this is known as an index equation) then just apply the method I showed you and you will be able to solve in future!

OpenStudy (anonymous):

Ok thanks :)